# Structured Probabilistic Inference & Generative Modeling workshop of NeurIPS 2025

### v2

@@ 3.1.2 Divergence-regularized sampling
where $D_{f}(q\|p)=\sum_{i}p(i)f\left(\frac{q(i)}{p(i)}\right)$ is an $f$-divergence with generator function $f:\mathbb{R}_{+}\to\mathbb{R}$, $q_{y}(\cdot)$ is a target distribution (discussed further in Appendix F.3), and $\alpha>0$ is the divergence weight. Different choices of $f$ lead to different divergences with distinct theoretical properties and practical implications for guidance.
@@
The gradient of the $f$-divergence regularized score can be written as:
$\nabla_{x}\mathcal{S}_{\mathcal{D}}(x,y)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)-\alpha\sum_{i=1}^{N}w_{f}(q_{y}(i),p(i|x))\cdot\nabla_{x}\log\frac{p(i|x)}{\bar{p}(x)},$ (8)
where $w_{f}(q,p)=p\cdot f^{\prime}\left(\frac{q}{p}\right)$ is the divergence-specific weighting function, $\nabla_{x}\log\frac{p(i|x)}{\bar{p}(x)}=\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)$, and $\bar{p}(x)$ denotes the normalizing constant.
@@ 3.1.3 Reverse KL Divergence: Mode-Covering Guidance
For reverse Kullback-Leibler divergence with $f(t)=-\log(t)$, we have $f^{\prime}(t)=-1/t$, which gives $w_{f}(q,p)=q$. This yields a particularly interpretable gradient form.
@@
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)+\alpha\sum_{i=1}^{N}q_{y}(i)\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$ (9)
@@
The reverse KL gradient can be decomposed as:
$\nabla_{x}D_{\text{KL}}(q_{y}\|p(\cdot|x))=\underbrace{\sum_{i=1}^{N}q_{y}(i)\nabla_{x}f_{i}(x)}_{\text{Target direction}}-\underbrace{\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)}_{\text{Current direction}}.$ (10)
@@
Let $X\sim\mathcal{P}$ be a random variable defined on $\mathbb{R}^{d}$, with density function $f(x)=\sum_{k=1}^{K}b_{k}f_{k}(x)$, where $f_{k}(x)$ is a normal density with mean $\mu_{k}$ and covariance matrix $\Sigma_{k}$, and $b_{k}>0$ with $\sum_{k=1}^{K}b_{k}=1$. Then:
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)=\sum_{k=1}^{K}\Sigma_{k}^{-1}(\mu_{k}-x)\cdot\Gamma_{k}(x,y,\alpha,\epsilon),$ (11)
@@
$\Gamma_{k}(x,y,\alpha,\epsilon)=\tau_{1}w_{k}(x)\mathbb{I}_{k=y}-\tau_{2}\sum_{j=1}^{K}w_{j}(x)p_{\tau_{2}}(k|x)+\alpha\left(q_{y}(k)-\sum_{j=1}^{K}w_{j}(x)p(k|x)\right),$ (12)
and $w_{k}(x)=\frac{b_{k}\exp\left(-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})\right)}{\sum_{j}b_{j}\exp\left(-\frac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\right)}$.
@@
When all covariances are identity matrices ($\Sigma_{k}=\sigma^{2}I$) and the target distribution follows $q_{y}(i)=(1-\epsilon)\frac{1}{N}+\epsilon\mathbb{I}_{i=y}$, the guidance becomes:
@@ 3.1.4 Forward KL Divergence: Mode-Seeking Guidance
In contrast to reverse KL, forward KL divergence with $f(t)=t\log(t)$ exhibits mode-seeking behavior, penalizing $p(\cdot|x)$ for placing mass where $q_{y}$ has little support. For this divergence, $f^{\prime}(t)=\log(t)+1$, giving $w_{f}(q,p)=q[\log(q/p)+1]$.
@@
$\nabla_{x}\mathcal{S}_{\text{FKL}}(x,y)$ $=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$ (14)
$\quad-\alpha\sum_{i=1}^{N}\left[q_{y}(i)\log\frac{q_{y}(i)}{p(i|x)}+q_{y}(i)\right]\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$
@@ 3.1.5 Jensen-Shannon Divergence: Balanced Guidance
$\nabla_{x}\mathcal{S}_{\text{JS}}(x,y)$ $=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$ (16)
$\quad-\frac{\alpha}{2}\sum_{i=1}^{N}\left[\frac{q_{y}(i)}{m(i)}-\frac{p(i|x)}{m(i)}\right]\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$
@@ C.3 Introduction to $f$-divergences
$f$-divergences [csiszar1967information, ali1966general] provide a general framework for measuring dissimilarity between probability distributions. For two probability distributions $P$ and $Q$ with densities $p$ and $q$ on a measurable space $\mathcal{X}$, an $f$-divergence is defined as:
$D_{f}(P\|Q)=\int_{\mathcal{X}}q(x)f\left(\frac{p(x)}{q(x)}\right)dx,$ (45)
where $f:\mathbb{R}_{+}\to\mathbb{R}$ is a convex function satisfying $f(1)=0$. The convexity of $f$ ensures $D_{f}(P\|Q)\geq 0$ with equality if and only if $P=Q$ almost everywhere.
For discrete distributions over $N$ outcomes, this simplifies to:
$D_{f}(p\|q)=\sum_{i=1}^{N}q(i)f\left(\frac{p(i)}{q(i)}\right).$ (46)
Different choices of $f$ recover classical divergences. Table 4 summarizes the four divergences used in this work, along with their generator functions and key properties.
Table 4: $f$-divergences used in this work with their generator functions and properties.
Divergence $f(t)$ $f^{\prime}(t)$ Key Property
Reverse KL $\log(t)$ 1/t-1/t Mode-covering, zero-avoiding
Forward KL $t\log(t)$ $\log(t)+1$ Mode-seeking, zero-forcing
Jensen-Shannon $(t+1)\log\frac{t+1}{2}+t\log t$ $\log\frac{t}{(t+1)/2}$ Symmetric, bounded
Squared Hellinger $(\sqrt{t}-1)^{2}$ $\frac{\sqrt{t}-1}{\sqrt{t}}$ Metric, moderate penalties
@@
For classifier-guided diffusion, we need $\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))$ where $p(\cdot|x)=\text{softmax}(f_{1}(x),\ldots,f_{N}(x))$ is the classifier output and $q_{y}$ is a fixed target distribution. Using the chain rule:
$\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}q(i)f^{\prime}\left(\frac{q_{y}(i)}{p(i|x)}\right)\frac{\partial}{\partial x}\left[\frac{q_{y}(i)}{p(i|x)}\right],$ (47)
which simplifies to Proposition 2 after applying the softmax gradient identity $\nabla_{x}\log p(i|x)=\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)$.
@@
The shape of $f$ determines the divergence’s behavior:
  • Asymmetry: Most $f$-divergences are asymmetric ($D_{f}(P\|Q)\neq D_{f}(Q\|P)$), except JS and Hellinger.
  • Tail behavior: When $t\to 0$ (i.e., $p(i)\to 0$ while $q_{y}(i)>0$), reverse KL has $f(t)\to\infty$ (zero-avoiding), while forward KL has $f(t)\to 0$ (zero-forcing).
  • Convexity: All $f$ are convex, ensuring the divergence is non-negative and has well-defined gradients.
The divergence choice fundamentally affects the guidance gradient’s weighting function $w_{f}(q,p)=p\cdot f^{\prime}(q/p)$, which determines how strongly each class contributes to the regularization term. For reverse KL, $w_{f}=q$ (linear in target); for forward KL, $w_{f}=q\log(q/p)+q$ (logarithmically penalizes excess mass); for Hellinger, $w_{f}=\sqrt{pq}$ (geometric mean damping).
@@ C.4 Squared Hellinger Divergence
For completeness, we also evaluated squared Hellinger distance, defined by $f(t)=(\sqrt{t}-1)^{2}$, which gives $f^{\prime}(t)=\frac{\sqrt{t}-1}{\sqrt{t}}=1-\frac{1}{\sqrt{t}}$ and $w_{f}(q,p)=p(1-1/\sqrt{q/p})=p-\sqrt{pq}$.
@@ D.2 Proof of Proposition 2
For any $f$-divergence $D_{f}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}p(i|x)f\left(\frac{q_{y}(i)}{p(i|x)}\right)$, taking the gradient with respect to $x$:
$\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))$ $=\nabla_{x}\sum_{i=1}^{N}p(i|x)f\left(\frac{q_{y}(i)}{p(i|x)}\right)$ (57)
$=\sum_{i=1}^{N}\left[\nabla_{x}p(i|x)\cdot f\left(\frac{q_{y}(i)}{p(i|x)}\right)+p(i|x)f^{\prime}\left(\frac{q_{y}(i)}{p(i|x)}\right)\nabla_{x}\frac{q_{y}(i)}{p(i|x)}\right].$ (58)
Since $q_{y}$ is independent of $x$, we have:
$\nabla_{x}\frac{q_{y}(i)}{p(i|x)}=-\frac{q_{y}(i)}{p(i|x)^{2}}\nabla_{x}p(i|x)=-\frac{q_{y}(i)}{p(i|x)^{2}}p(i|x)\nabla_{x}\log p(i|x)=-\frac{q_{y}(i)}{p(i|x)}\nabla_{x}\log p(i|x).$ (59)
Substituting back:
$\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))$ $=\sum_{i=1}^{N}\left[p(i|x)\nabla_{x}\log p(i|x)\cdot f\left(\frac{q_{y}(i)}{p(i|x)}\right)-p(i|x)f^{\prime}\left(\frac{q_{y}(i)}{p(i|x)}\right)\frac{q_{y}(i)}{p(i|x)}\nabla_{x}\log p(i|x)\right]$ (60)
$=\sum_{i=1}^{N}p(i|x)\nabla_{x}\log p(i|x)\left[f\left(\frac{q_{y}(i)}{p(i|x)}\right)-f^{\prime}\left(\frac{q_{y}(i)}{p(i|x)}\right)\frac{q_{y}(i)}{p(i|x)}\right].$ (61)
Using $\nabla_{x}\log p(i|x)=\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)$ and defining $w_{f}(q,p)=p\cdot f^{\prime}(q/p)$, we obtain the stated result. Combined with $\nabla_{x}\log p_{\tau_{1},\tau_{2}}(y|x)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$ from Equation (7), the proposition follows. ∎

### v3

@@ 3.1.2 Divergence-regularized sampling
where $D_{f}(q\|p)=\sum_{i}q(i)f\left(\frac{p(i)}{q(i)}\right)$ is an $f$-divergence with generator function $f:\mathbb{R}_{+}\to\mathbb{R}$, $q_{y}(\cdot)$ is a target distribution (discussed further in Appendix F.3), and $\alpha>0$ is the divergence weight. We use $q_{y}(i)=(1-\epsilon)\frac{1}{N}+\epsilon\mathbb{I}_{i=y}$ with $\epsilon=0.1$, where the uniform component $(1-\epsilon)/N>0$ ensures $q_{y}(i)>0$ for all $i$. This smoothing is essential: if $q_{y}$ were one-hot, the divergence $D_{f}(q_{y}\|p)$ would be infinite unless $p$ is also one-hot (which never occurs during sampling). The uniform component prevents divergence blow-up while $\epsilon$ provides target class emphasis. Different choices of $f$ lead to different divergences with distinct theoretical properties and practical implications for guidance.
Assumption 1.
The generator function $f:\mathbb{R}_{+}\to\mathbb{R}$ satisfies:
  1. 1.
    $f$ is differentiable on $(0,\infty)$,
  2. 2.
    $f(1)=0$,
  3. 3.
    The target distribution satisfies $q_{y}(i)>0$ for all $i$ to ensure finite divergence.
@@
The gradient of the $f$-divergence regularized score is
$\nabla_{x}\mathcal{S}_{\mathcal{D}}(x,y)$ $=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\!\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\,\nabla_{x}f_{i}(x)$
$\quad-\alpha\!\sum_{i=1}^{N}w_{f}(q_{y}(i),p(i|x))\,g_{i}(x),$ (8)
where $p(i|x)=\frac{\exp(f_{i}(x))}{\sum_{j}\exp(f_{j}(x))}$, $g_{i}(x)=\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\,\nabla_{x}f_{j}(x)$, $p_{\tau_{2}}(i|x)=\frac{\exp(\tau_{2}f_{i}(x))}{\sum_{j}\exp(\tau_{2}f_{j}(x))}$ and $w_{f}(q,p)=p\,f^{\prime}\!\left(\frac{p}{q}\right)$.
@@ 3.1.3 Reverse KL Divergence: Mode-Covering Guidance
For reverse Kullback-Leibler divergence with $f(t)=-\log(t)$, we have $f^{\prime}(t)=-1/t$, which gives $w_{f}(q,p)=-q$. This yields a particularly interpretable gradient form.
@@
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)=$ $\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$ (9)
$+\alpha\sum_{i=1}^{N}q_{y}(i)\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$
@@
The negative of reverse KL gradient can be decomposed as:
[…]
@@
Let $X\sim\mathcal{P}$ be a random variable defined on $\mathbb{R}^{d}$, with density function $f(x)=\sum_{k=1}^{K}b_{k}f_{k}(x)$, where $f_{k}(x)$ is a normal density with mean $\mu_{k}$ and covariance matrix $\Sigma_{k}$, and $b_{k}>0$ with $\sum_{k=1}^{K}b_{k}=1$. Define the posterior as $w_{k}(x):=p(k|x)=\frac{b_{k}f_{k}(x)}{\sum_{j=1}^{K}b_{j}f_{j}(x)}$. Then:
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)=\sum_{k=1}^{K}f_{k}(x)\Sigma_{k}^{-1}(\mu_{k}-x)\,\Gamma_{k}(x,y,\alpha,\epsilon),$ (11)
@@
$\Gamma_{k}(x,y,\alpha,\epsilon)=\tau_{1}\mathbb{I}_{k=y}-\tau_{2}\,p_{\tau_{2}}(k|x)+\alpha\big(q_{y}(k)-w_{k}(x)\big),$ (12)
and $w_{k}(x)$ is the posterior distribution. Here $p_{\tau_{2}}(k|x)$ is the tempered posterior over components, e.g. $p_{\tau_{2}}(k|x)\propto\exp(\tau_{2}\,\ell_{k}(x))$ with $\ell_{k}(x)=\log b_{k}+\log f_{k}(x)$, and $\epsilon$ enters only through $q_{y}$ when you choose a smoothed target.
@@
When all covariances are identity matrices ($\Sigma_{k}=\sigma^{2}I$) and the target distribution follows $q_{y}(i)=(1-\epsilon)\frac{1}{N}+\epsilon\mathbb{I}_{i=y}$, the guidance is proportional to:
@@ 3.1.4 Forward KL Divergence: Mode-Seeking Guidance
In contrast to reverse KL, forward KL divergence with $f(t)=t\log(t)$ exhibits mode-seeking behavior, penalizing $p(\cdot|x)$ for placing mass where $q_{y}(\cdot)$ has little support. For this divergence, we have $w_{f}(q,p)=p[\log(p/q)+1]$.
@@
$\nabla_{x}S_{FKL}(x,y)$ $=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$
$\quad-\alpha\sum_{i=1}^{N}p(i|x)\left[\log\frac{p(i|x)}{q_{y}(i)}+1\right]$
$\quad\quad\times\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$
@@ 3.1.5 Jensen-Shannon Divergence: Balanced Guidance
$\nabla_{x}S_{JS}(x,y)$ $=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)$
$\quad-\alpha\sum_{i=1}^{N}p(i|x)\log\frac{2p(i|x)}{q_{y}(i)+p(i|x)}$
$\quad\quad\times\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$
where $g_{i}(x)=\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)$.
@@ C.3 Introduction to $f$-divergences
$f$-divergences [csiszar1967information, ali1966general] provide a general framework for measuring dissimilarity between probability distributions. In this work we adopt the discrete convention
$D_{f}(q\|p)\;=\;\sum_{i=1}^{N}q(i)\,f\!\left(\frac{p(i)}{q(i)}\right),$ (43)
where $f:\mathbb{R}_{+}\to\mathbb{R}$ is convex and satisfies $f(1)=0$. Convexity implies $D_{f}(q\|p)\geq 0$ with equality iff $p=q$.
Different choices of $f$ recover classical divergences; Table 4 summarizes the divergences used.
Table 4: $f$-divergences used in this work with their generator functions.
Divergence $f(t)$ $f^{\prime}(t)$ Notes
Reverse KL ($D_{\mathrm{KL}}(q\|p)$) $\log t$ 1/t-1/t penalizes $p\!\to\!0$ when q>0q\!>\!0
Forward KL ($D_{\mathrm{KL}}(p\|q)$) $t\log t$ $\log t+1$ penalizes p>0p\!>\!0 when $q\!\to\!0$
Jensen–Shannon $(t+1)\log\frac{t+1}{2}+t\log t$ $\log\frac{2t}{t+1}$ symmetric, bounded
Squared Hellinger $(\sqrt{t}-1)^{2}$ $1-\frac{1}{\sqrt{t}}$ $H$ is a metric (not $H^{2}$)
@@
Let $p(\cdot|x)$ denote the classifier output over NN classes and $q_{y}$ a fixed target distribution. For (43), define
$w_{f}(q,p)\;:=\;p\,f^{\prime}\!\left(\frac{p}{q}\right).$ (44)
Using $\nabla_{x}p(i|x)=p(i|x)\nabla_{x}\log p(i|x)$, we obtain
$\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))\;=\;\sum_{i=1}^{N}w_{f}\big(q_{y}(i),p(i|x)\big)\,\nabla_{x}\log p(i|x),$ (45)
and with the softmax identity $\nabla_{x}\log p(i|x)=\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x),$ this yields Proposition 2.
@@
Most $f$-divergences are asymmetric, while JS and Hellinger are symmetric. The choice of $f$ determines the weighting $w_{f}(q_{y}(i),p(i|x))$ in (45): for reverse KL, $w_{f}(q,p)=-q$ (non-vanishing as $p\to 0$); for forward KL, $w_{f}(q,p)=p\big(\log\frac{p}{q}+1\big)$ (vanishes as $p\to 0$); for squared Hellinger, $w_{f}(q,p)=p-\sqrt{pq}$ (damped by the geometric mean term). These weightings help explain the different empirical precision–recall behaviors observed in Section 4.
@@ C.4 Squared Hellinger Divergence
For completeness, we also evaluated squared Hellinger distance, defined by $f(t)=(\sqrt{t}-1)^{2}$, with derivative $f^{\prime}(t)=1-\frac{1}{\sqrt{t}}$. Under our convention $t=\frac{p}{q}$, the corresponding weight is $w_{f}(q,p)=pf^{\prime}(p/q)=p-\sqrt{pq}$.
@@ D.2 Proof of Proposition 2
Let $D_{f}(q_{y}\|p)=\sum_{i}q_{i}\,f\!\left(\frac{p_{i}}{q_{i}}\right),$ where $p_{i}:=p(i|x),\;q_{i}:=q_{y}(i).$ Taking the gradient with respect to $x$, and noting that $q_{i}$ is constant:
$\nabla_{x}D_{f}$ $=\sum_{i}q_{i}\,f^{\prime}\!\left(\frac{p_{i}}{q_{i}}\right)\nabla_{x}\!\left(\frac{p_{i}}{q_{i}}\right)$
$=\sum_{i}f^{\prime}\!\left(\frac{p_{i}}{q_{i}}\right)\nabla_{x}p_{i}.$ (55)
Using $\nabla_{x}p_{i}=p_{i}\,\nabla_{x}\log p_{i}$ and $\nabla_{x}\log p_{i}=\nabla_{x}f_{i}(x)-\sum_{j}p_{j}\,\nabla_{x}f_{j}(x)=:g_{i}(x),$ we obtain
$\nabla_{x}D_{f}=\sum_{i}p_{i}\,f^{\prime}\!\left(\frac{p_{i}}{q_{i}}\right)g_{i}(x)=\sum_{i}w_{f}(q_{i},p_{i})\,g_{i}(x).$ (56)
Combining with $\nabla_{x}\log p_{\tau_{1},\tau_{2}}(y|x)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i}p_{\tau_{2}}(i|x)\,\nabla_{x}f_{i}(x)$ and the negative sign in $\mathcal{S}_{\mathcal{D}}=\log p_{\tau_{1},\tau_{2}}(y|x)-\alpha D_{f}(q_{y}\|p)$ gives Proposition 2. ∎